Without any taxes, equity has a return of 24%, and debt has a return of 12%. The overall WACC is 16%. What is the debt ratio (i.e., the proportion of debt financing)?

66.7%

75.0%

60.0%

or 54.0%
Answer #1

WACC = Wd×Rd×(1-t)+We×Ke

W is weights of respective portfolios

R is return on respective portfolios

Wd+We = 1

16% = Wd×12%×(1-0%)+(1-Wd)×24%

16% = Wd×12%+24%-Wd×24%

Wd×12% = 8%

Hence, Debt ratio is 66.7%

*Please rate thumbs up